# Which symmetry group for elementary particles with an electric charge today and in the past?

### v1

@@ 5 Elementary particles in Kaluza-Klein relativity
where the scalar $q$ and $Q\in\mathbb{R}^{4}$ are extra momenta with respect to those of Poincaré’s group. The natural conjecture is to identify $q$ to the electric charge with suitable physical units. This idea is mentioned in Petit and D’Agostini [2014] but the extension of Poincaré’s group proposed by the authors is only a subgroup of $\hat{\mathbb{G}}_{1}$ of dimension 11. In the present work, the aim is to perform an in-depth investigation by examining the consequences of this conjecture to verify if they fit the observations.
@@
6 Elementary particles in an Universe with a microscopic fifth dimension
@@ 6.1 Zoom in
To avert the previous paradox concerning the electric charge, it is worth to remark that the fifth dimension is curled up and the estimate of the cylinder radius (about $10^{-30}$ cm) is overwhelmingly small with respect to usual lengths in nuclear physics. The idea is to take advantage of this fact by a zoom in along the fifth coordinate and to consider the limit when the cylinder radius $\omega$ vanishes. Then we consider the transformation law of a vector $\hat{\bm{V}}$
@@
The first matrix represents a covariant semi-metric of signature (+0)(+---0)( + - - - 0 ) while the second one represents a contravariant semi-metric of signature (0 0 0 0)limit-from0000(0\,0\,0\,0-)( 0 0 0 0 - ). After scaling, the Euclidean structure of $\hat{\mathcal{U}}$ is canceled but it remains two debris, a symmetric 2-covariant tensor that we denote $\hat{\bm{G}_{0}}$ represented by $\hat{G}_{0}$ and a symmetric 2-contravariant tensor or, equivalently, a vector $\bm{\Omega}_{0}$ represented by the column
@@
The set $\hat{\mathbb{G}}_{0}$ of affine transformations $\hat{a}=(\hat{C},\hat{P})$ of $\mathbb{R}^{5}$ of which the linear part conserves the components of $\hat{\bm{G}_{0}}$ and $\bm{\Omega}_{0}$ is such that
@@ 6.2 Gauge transformation
To understand the meaning of $b$, we change of point of view and work in the framework of the general relativity in which $\hat{\bm{V}}$ is a tangent vector to a the 5D manifold $\hat{\mathcal{U}}$. As in de Saxcé and Vallée [2016], we hope to determine whether the G-structure in Kobayashi sense (Kobayashi [1972]) is integrable by solving thanks to Frobenius method the PDE system
@@
The $\hat{\mathbb{G}}_{0}$-structure is integrable only if the manifold is flat. Otherwise, it is convenient as in Souriau [1964, Chapter VII, (41.21)] to consider the transition maps of standard charts
@@ 6.3 Coadjoint orbit
$M=P\,M^{\prime}P^{*}+C\,(P\,(\Pi^{\prime}-q\,b))^{*}-(P\,(\Pi^{\prime}-q\,b))\,\hat{C}^{*}+(P\,b)(P\,Q^{\prime})^{*}-(P\,Q^{\prime})(P\,b)^{*}$ (41)
@@
$M=P\,M^{\prime}P^{*}+C\,\Pi^{*}-\Pi\,\hat{C}^{*}+(P\,b)(P\,Q^{\prime})^{*}-(P\,Q^{\prime})(P\,b)^{*}$ (45)
@@
To author’s knowledge, the Physics of the symmetry group $\hat{\mathbb{G}}_{0}$ has not been considered in the literature. By the way, there exists a group introduced by Lévy-Leblond [1965] and called by him Carroll’s group [^4] Lewis Carroll, the author of Alice’s Adventures in Wonderland, although mathematician, had nothing to do with the creation of this group., of which the elements have the same mathematical structure as the ones of $\hat{\mathbb{G}}_{0}$. However the physical meaning of Carroll’s group is very different from the one of $\hat{\mathbb{G}}_{0}$: Carroll’s group linearly acts on a space of dimension 4 (not 5) and the dimension analogous to the fifth of $\hat{\mathbb{G}}_{0}$ is timelike (not spacelike). Carroll’s group was constructed as a degeneracy of Poincaré’s group (similar but distinct from Galileo’s group). Despite of these discrepancies concerning the physical aspects, Carroll’s group may be considered as a forunner of the group $\hat{\mathbb{G}}_{0}$. For more details on Carroll’s group, the reader is referred to Duval et al. [1985, 1991, 2014], Bergshoeff et al. [2014, 2023], Lévy-Leblond [2023].
@@ 7 A cosmological scenario for the evolution of elementary particle structure
One of the important problems in Kaluza-Klein theories is how to explain the large separation of the scale of our 3D space and that of the extra dimension. This issue was addressed in Okada [1986] with a Robertson-Walker metric in our four dimensions and a $d$-dimensional sphere $S^{d}$ in the extra dimensions. For our concern, $d=1$. Under certain assumptions, this approach predicts stable cosmological models for the extra dimension and, using Kasner solutions for approximations when $t\rightarrow 0$, a possible evolution of the scale factors $a$ for our 3D space and $\omega$ for the fifth dimension given by
@@
  • next the three former space dimensions inflate quickly while the fifth one shrinks,
@@
When $\omega=0$, we recover the group $\hat{\mathbb{G}}_{0}$. We would like to classify the elementary particles with charge at scale $\omega>0$. In particular, we shall obtain the classification for the early Universe ($\omega=1$) of which the study was only sketched out in Section 5. The other limit case of the Universe today ($\omega=0)$ that is singular was already studied in Section 6.
@@ 7.1 Polarization plane and map
At this stage, it is worth to remark an important difference between Poincaré relativity, in which $A_{M_{0}}$ and its adjoint *$*A_{M_{0}}$ are both 2-forms, and Kaluza-Klein one for which $A_{\hat{M}_{0}}$ is a 3-form while *$*A_{\hat{M}_{0}}$ is a 2-form, that has important consequences from the geometrical and physical viewpoints. Owing to (60), $\hat{M}_{0}$ does not depends on $\hat{X}$ when $\hat{X}$ varies over $\hat{\mathcal{D}}$. Let $\hat{I},\hat{J}_{1},\hat{J}_{2}\in\mathbb{R}^{5}$ such that
@@
*$*A_{\hat{M}}(\hat{\Pi},\hat{J}_{k})=*A_{\hat{M}_{0}}(\hat{\Pi},\hat{J}_{k})+vol(\hat{\Pi},\hat{X},\hat{\Pi},\hat{J}_{k})=s\,*A_{\hat{\Omega}}(\hat{\Pi},\hat{J}_{k})\qquad(k=1,2)$
that leads to investigate the properties of the 2-form *$*A_{\hat{\Omega}}$. Applying Theorem 2.1 $\heartsuit$ and $\spadesuit$, its adjoint is the 3-form
*$**A_{\hat{\Omega}}=-*vol(\hat{I},\hat{J}_{1},\hat{J}_{2})=-\hat{I}^{*}\wedge\hat{J}^{*}_{1}\wedge\hat{J}^{*}_{2}$
@@
$*\hat{J}_{1}=-(*A_{\hat{\Omega}}(\hat{I},\hat{J}_{2}))^{*},\qquad\hat{J}_{2}=(*A_{\hat{\Omega}}(\hat{I},\hat{J}_{1}))^{*}$ (63)
@@
$*\mbox{pol}_{\hat{\mu}}:\mathbb{R}^{5}\rightarrow\mathbb{R}^{5}:\hat{U}\mapsto\hat{V}=\mbox{pol}_{\hat{\mu}}(\hat{U})=(*A_{\hat{M}}(\hat{\Pi},\hat{U}))^{*}$
@@
    then $\hat{J}^{\prime}_{1}$ and $\hat{J}^{\prime}_{2}$ are linear combinations of $\hat{J}_{1}$ and $\hat{J}_{2}$. This ensures the uniqueness of the polarization plane. In contrast, the orthonormal basis $(\hat{J}_{1},\hat{J}_{2})$ is defined only modulo an orthogonal transformation of $\mathcal{P}$
@@ 7.2 Invariants of the motion
$m_{0}=\sqrt{\hat{\Pi}^{*}\hat{\Pi}},\qquad s=\frac{\sqrt{-(\mbox{pol}_{\hat{\mu}}(\hat{J}_{1}))^{*}\mbox{pol}_{\hat{\mu}}(\hat{J}_{1})}}{\sqrt{\hat{\Pi}^{*}\hat{\Pi}}\,\sqrt{-(\hat{J}_{1})^{*}\hat{J}_{1}}}$
@@ 8 Constructing the connection associated to the group $\hat{\mathbb{G}}_{0}$
At every symmetry group $G$ is attached a family of covariant derivatives of which the connection matrices are defined on the $G$-principal bundle of the $G$-frames and valued in its Lie algebra $\mathfrak{g}$ (see for instance Kobayashi and Nomizu [1963]). We call them $G$-connections. For the group $\hat{\mathbb{G}}_{1}$, we recover the Levi-Civita free torsion connection used in Kaluza-Klein theory. A key-point of interest is to determine the $\hat{\mathbb{G}}_{0}$-connections, the difficulty being that the manifold $\hat{\mathcal{U}}$ is not Riemannian, and to use it to deduce the equation of motion of a particle in gravitation and electromagnetic fields.
@@ 8.1 Zoom out
It is time now to restore the primes canceled for convenience in Section 6.1 after the zoom in. For instance, according to (35) and (36), the two tensors $\hat{\bm{G}_{0}}$ and $\bm{\Omega}_{0}$ are represented at the microscopic scale respectively by
@@ 8.2 Hypotheses on the pullback connection
In order to avert the singularity resulting from the zoom out without to lose the relevant information in the fifth dimension, we propose to pullback over the space-time $\mathcal{U}$ the $\hat{\mathbb{G}}_{0}$-connection $\hat{\nabla}$ on the tangent bundle $T\hat{\mathcal{U}}$. They are an infinite number of $\hat{\mathbb{G}}_{0}$-connections. The purpose here is not to study their properties in generic terms and to classify them. We are interested only by those which satisfy certain physical requirements on the ground of hypotheses considered below, (H1) to (H4) on the underlying geometric structure of the Universe and (H5) on the motion of a charged elementary particle in both the gravitation and electromagnetic fields.
@@
    is a field of symmetric 2-covariant tensors (but not a metric!). $(\hat{\bm{e}}^{\prime}_{\alpha})$ is a $\hat{\mathbb{G}}_{0}$-basis, i.e. a basis in which $\hat{\bm{G}_{0}}$ and $\bm{\Omega}_{0}$ are represented by (68).
@@
    $\hat{\nabla}$ is a $\hat{\mathbb{G}}_{0}$-connection. Let $\pi:\mathcal{F}(\hat{\mathcal{U}})\rightarrow\hat{\mathcal{U}}$ be the frame bundle of $\hat{\mathcal{U}}$, i.e. the bundle whose fiber over $\bm{\hat{X}}$ is the set of all basis of $T_{\bm{\hat{X}}}\hat{\mathcal{U}}$. A moving frame is a smooth section $\bm{\hat{X}}\mapsto(\hat{\bm{e}}^{\prime}_{\alpha})$ of $\mathcal{F}$. It is called a coordinate frame (or natural frame) if locally the basis is associated to a coordinate system (or local chart) $(\hat{X}^{\alpha})$ and if so it is denoted (^$(\hat{\bm{\partial}}_{\alpha})$. The set $\mathcal{F}_{\hat{\bm{G}_{0}}}$ of $\hat{\mathbb{G}}_{0}$-bases is called $\hat{\mathbb{G}}_{0}$-structure. It is a subbundle of $\mathcal{F}(\hat{\mathcal{U}})$ and a principal $\hat{\mathbb{G}}_{0}$-bundle that is endowed with an Ehresmann connection. A $\hat{\mathbb{G}}_{0}$-connection $\hat{\nabla}$ is defined by the field of $\mathfrak{g}_{0}$-valued connection 1-form $(\hat{\bm{e}}^{\prime}_{\alpha})\mapsto\hat{\Gamma}^{\prime}$ on $\mathcal{F}_{\hat{\bm{G}_{0}}}$ such that
@@
    The section of the fiber bundle $\pi:\hat{\mathcal{U}}\rightarrow\mathcal{U}$ is an embedding of $\mathcal{U}$ into $\hat{\mathcal{U}}$. The subspace $H_{f(\bm{X})}=(Tf)T_{\bm{X}}\mathcal{U}$ of $T_{f(\bm{X})}\hat{\mathcal{U}}$ is isomorphic to $T_{\bm{X}}\mathcal{U}$ then of dimension 4 and the tangent space to the fiber $\hat{\mathcal{U}}_{\bm{X}}$ at $f(\bm{X})$ is the kernel of $T\pi$. Then $T_{f(\bm{X})}\hat{\mathcal{U}}=H_{f(\bm{X})}\oplus T_{f(\bm{X})}\hat{\mathcal{U}}_{\bm{X}}$. A basis adapted to the section $f$ is a fibered basis $(\hat{\bm{e}}_{1},\cdot,\hat{\bm{e}}_{5})$ such that $(\hat{\bm{e}}_{1},\cdot,\hat{\bm{e}}_{4})$ is a basis of $H_{f(\bm{X})}$. A way to build such a basis is to choose a non null vector $\hat{\bm{e}}_{5}$ tangent to the fiber $\hat{\mathcal{U}}_{\bm{X}}$ and a basis $(\bm{e}_{1},\cdot,\bm{e}_{4})$ of $T_{\bm{X}}\mathcal{U}$. Then, with the convention that Greek indices run from 1 to 5 and Latin ones run from 1 to 4, $(\hat{\bm{e}}_{1},\cdot,\hat{\bm{e}}_{4},\hat{\bm{e}}_{5})=((Tf)\bm{e}_{1},\cdot,(Tf)\hat{\bm{e}}_{4},\hat{\bm{e}}_{5})$ is a basis adapted to the section $f$ and
@@
    In terms of the dual bases, on has

### v2

@@ 5 Elementary particles in Kaluza-Klein relativity
where the scalar $q$ and $Q\in\mathbb{R}^{4}$ are extra momenta with respect to those of Poincaré’s group. The natural conjecture is to identify $q$ to the electric charge with suitable physical units. This idea is mentioned in [39] but the extension of Poincaré’s group proposed by the authors is only a subgroup of $\hat{\mathbb{G}}_{1}$ of dimension 11. In the present work, the aim is to perform an in-depth investigation by examining the consequences of this conjecture to verify if they fit the observations.
@@
6 Elementary particles in an Universe with an overwhelmingly small fifth dimension
@@ 6.1 Zoom in
To avert the previous paradox concerning the electric charge, it is worth to remark that the fifth dimension is curled up and the estimate of the cylinder radius (about $10^{-31}$ cm) is overwhelmingly small with respect to usual lengths in nuclear physics. The idea is to take advantage of this fact by a zoom in along the fifth coordinate and to consider the limit when the cylinder radius $\omega$ vanishes. Then we consider the transformation law of a vector $\hat{\bm{V}}$
@@
The first matrix represents a covariant semi-metric of signature (+0)(+---0)( + - - - 0 ) while the second one represents a contravariant semi-metric of signature (0 0 0 0)limit-from0000(0\,0\,0\,0-)( 0 0 0 0 - ). After scaling, the Euclidean structure of $\hat{\mathcal{U}}$ is canceled but it remains two debris, a symmetric 2-covariant tensor that we denote $\hat{\bm{G}_{0}}$ represented by $\hat{G}_{0}$ and a symmetric 2-contravariant tensor or, equivalently, a vector $\hat{\bm{\Omega}}_{0}$ represented by the column
@@
The set $\hat{\mathbb{G}}_{0}$ of affine transformations $\hat{a}=(\hat{C},\hat{P})$ of $\mathbb{R}^{5}$ of which the linear part conserves the components of $\hat{\bm{G}_{0}}$ and $\hat{\bm{\Omega}}_{0}$ is such that
@@ 6.2 Gauge transformation
To understand the meaning of $b$, we change of point of view and work in the framework of the general relativity in which $\hat{\bm{V}}$ is a tangent vector to a the 5D manifold $\hat{\mathcal{U}}$. As in [10], we hope to determine whether the G-structure in Kobayashi sense ([31]) is integrable by solving thanks to Frobenius method the PDE system
@@
The $\hat{\mathbb{G}}_{0}$-structure is integrable only if the manifold is flat. Otherwise, it is convenient as in Chapter VII, (41.21) of [45] to consider the transition maps of standard charts
@@ 6.3 Coadjoint orbit
$M=P\,M^{\prime}P^{*}+C\,(P\,(\Pi^{\prime}-q\,b))^{*}-(P\,(\Pi^{\prime}-q\,b))\,C^{*}+(P\,b)(P\,Q^{\prime})^{*}-(P\,Q^{\prime})(P\,b)^{*}$ (41)
@@
$M=P\,M^{\prime}P^{*}+C\,\Pi^{*}-\Pi\,C^{*}+(P\,b)(P\,Q^{\prime})^{*}-(P\,Q^{\prime})(P\,b)^{*}$ (45)
@@
To author’s knowledge, the Physics of the symmetry group $\hat{\mathbb{G}}_{0}$ has not been considered in the literature. By the way, there exists a group introduced by [33] and called by him Carroll’s group [^4] Lewis Carroll, the author of Alice’s Adventures in Wonderland, although mathematician, had nothing to do with the creation of this group., of which the elements have the same mathematical structure as the ones of $\hat{\mathbb{G}}_{0}$. However the physical meaning of Carroll’s group is very different from the one of $\hat{\mathbb{G}}_{0}$: Carroll’s group linearly acts on a space of dimension 4 (not 5) and the dimension analogous to the fifth of $\hat{\mathbb{G}}_{0}$ is timelike (not spacelike). Carroll’s group was constructed as a degeneracy of Poincaré’s group (similar but distinct from Galileo’s group). Despite of these discrepancies concerning the physical aspects, Carroll’s group may be considered as a forunner of the group $\hat{\mathbb{G}}_{0}$. For more details on Carroll’s group, the reader is referred to [12, 13, 14, 3, 4, 34].
@@ 7 A cosmological scenario for the evolution of elementary particle structure
One of the important problems in Kaluza-Klein theories is how to explain the large separation of the scale of our 3D space and that of the extra dimension. This issue was addressed in [37] with a Robertson-Walker metric in our four dimensions and a $d$-dimensional sphere $S^{d}$ in the extra dimensions. For our concern, $d=1$. Under certain assumptions, this approach predicts stable cosmological models for the extra dimension and, using Kasner solutions for approximations when $t\rightarrow 0$, a possible evolution of the scale factors $a$ for our 3D space and $\omega$ for the fifth dimension given by
@@
  • next the three former space dimensions inflate quickly while the last one shrinks,
@@
We would like to classify the elementary particles with charge at scale $\omega>0$. In particular, we shall obtain the classification for the early Universe ($\omega=1$) of which the study was only sketched out in Section 5. The other limit case of the Universe today ($\omega=0)$ that is singular was already studied in Section 6.
@@ 7.1 Polarization plane and map
At this stage, it is worth to remark an important difference between Poincaré relativity, in which $A_{M_{0}}$ and its adjoint $*A_{M_{0}}$ are both 2-forms, and Kaluza-Klein one for which $A_{\hat{M}_{0}}$ is a 3-form while $*A_{\hat{M}_{0}}$ is a 2-form, that has important consequences from the geometrical and physical viewpoints. Owing to (60), $\hat{M}_{0}$ does not depends on $\hat{X}$ when $\hat{X}$ varies over $\hat{\mathcal{D}}$. Let $\hat{I},\hat{J}_{1},\hat{J}_{2}\in\mathbb{R}^{5}$ such that
@@
$*A_{\hat{M}}(\hat{\Pi},\hat{J}_{k})=*A_{\hat{M}_{0}}(\hat{\Pi},\hat{J}_{k})+vol(\hat{\Pi},\hat{X},\hat{\Pi},\hat{J}_{k})=s\,*A_{\hat{\Omega}}(\hat{\Pi},\hat{J}_{k})\qquad(k=1,2)$
that leads to investigate the properties of the 2-form $*A_{\hat{\Omega}}$. Applying Theorem 2.1 $\heartsuit$ and $\spadesuit$, its adjoint is the 3-form
$*A_{\hat{\Omega}}=-*vol(\hat{I},\hat{J}_{1},\hat{J}_{2})=-\hat{I}^{*}\wedge\hat{J}^{*}_{1}\wedge\hat{J}^{*}_{2}$
@@
$\hat{J}_{1}=-(*A_{\hat{\Omega}}(\hat{I},\hat{J}_{2}))^{*},\qquad\hat{J}_{2}=(*A_{\hat{\Omega}}(\hat{I},\hat{J}_{1}))^{*}$ (63)
@@
$\mbox{pol}_{\hat{\mu}}:\mathbb{R}^{5}\rightarrow\mathbb{R}^{5}:\hat{V}\mapsto\hat{J}=\mbox{pol}_{\hat{\mu}}(\hat{V})=(*A_{\hat{M}}(\hat{\Pi},\hat{V}))^{*}$
@@
    then $\hat{J}^{\prime}_{1}$ and $\hat{J}^{\prime}_{2}$ are linear combinations of $\hat{J}_{1}$ and $\hat{J}_{2}$. This ensures the uniqueness of the polarization plane. In contrast, the orthonormal basis $(\hat{J}_{1},\hat{J}_{2})$ is defined only modulo an orthogonal transformation of $\mathcal{P}$.
@@ 7.2 Invariants of the motion
$m_{0}=\sqrt{\hat{\Pi}^{*}\hat{\Pi}},\qquad s=\frac{\sqrt{-(\mbox{pol}_{\hat{\mu}}(\hat{J}_{1}))^{*}\mbox{pol}_{\hat{\mu}}(\hat{J}_{1})}}{\sqrt{\hat{\Pi}^{*}\hat{\Pi}}}$
@@ 8 Constructing the connection associated to the group $\hat{\mathbb{G}}_{0}$
At every symmetry group $G$ is attached a family of covariant derivatives of which the connection matrices are defined on the $G$-principal bundle of the $G$-frames and valued in its Lie algebra $\mathfrak{g}$ (see for instance [30]). We call them $G$-connections. For the group $\hat{\mathbb{G}}_{1}$, we recover the Levi-Civita free torsion connection used in Kaluza-Klein theory. A key-point of interest is to determine the $\hat{\mathbb{G}}_{0}$-connections, the difficulty being that the manifold $\hat{\mathcal{U}}$ is not Riemannian, and to use it to deduce the equation of motion of a particle in gravitation and electromagnetic fields.
@@ 8.1 Zoom out
It is time now to restore the primes canceled for convenience in Section 6.1 after the zoom in. For instance, according to (35) and (36), the two tensors $\hat{\bm{G}_{0}}$ and $\hat{\bm{\Omega}}_{0}$ are represented at the small scale respectively by
@@ 8.2 Hypotheses on the pullback connection
In order to avert the singularity resulting from the zoom out without to lose the relevant information in the fifth dimension, we propose to pullback over the space-time $\mathcal{U}$ the $\hat{\mathbb{G}}_{0}$-connection $\hat{\nabla}$ on the tangent bundle $T\hat{\mathcal{U}}$. The number of $\hat{\mathbb{G}}_{0}$-connections is infinite. The purpose here is not to study their properties in generic terms and to classify them. We are interested only by those which satisfy certain physical requirements on the ground of hypotheses considered below, (H1) to (H4) on the underlying geometric structure of the Universe and (H5) on the motion of a charged elementary particle in both the gravitation and electromagnetic fields.
@@
    is a field of symmetric 2-covariant tensors (but not a metric!). $(\hat{\bm{e}}^{\prime}_{\alpha})$ is a $\hat{\mathbb{G}}_{0}$-basis, i.e. a basis in which $\hat{\bm{G}_{0}}$ and $\hat{\bm{\Omega}}_{0}$ are represented by (68).
@@
    $\hat{\nabla}$ is a $\hat{\mathbb{G}}_{0}$-connection. Let $\hat{\pi}:\mathcal{F}(\hat{\mathcal{U}})\rightarrow\hat{\mathcal{U}}$ be the frame bundle of $\hat{\mathcal{U}}$, i.e. the bundle whose fiber over $\bm{\hat{X}}$ is the set of all basis of $T_{\bm{\hat{X}}}\hat{\mathcal{U}}$. A moving frame is a smooth section $\bm{\hat{X}}\mapsto(\hat{\bm{e}}^{\prime}_{\alpha})$ of $\mathcal{F}$. It is called a coordinate frame (or natural frame) if locally the basis is associated to a coordinate system (or local chart) $(\hat{X}^{\alpha})$ and if so it is denoted (^$(\hat{\bm{\partial}}_{\alpha})$. The set $\mathcal{F}_{\hat{\bm{G}_{0}}}$ of $\hat{\mathbb{G}}_{0}$-bases is called $\hat{\mathbb{G}}_{0}$-structure. It is a subbundle of $\mathcal{F}(\hat{\mathcal{U}})$ and a principal $\hat{\mathbb{G}}_{0}$-bundle that is endowed with an Ehresmann connection.
    As we are concerned only by the set of basis changes, i.e. the Lie subgroup of linear transformations $\hat{a}=(0,\hat{P})$ of $\hat{\bm{G}_{0}}$, the corresponding infinitesimal generators of $\hat{\mathfrak{g}}_{0}$ are of the form $\delta\hat{a}=(0,\delta\hat{P})$. A $\hat{\mathbb{G}}_{0}$-connection $\hat{\nabla}$ is defined by the field of $\hat{\mathfrak{g}}_{0}$-valued connection 1-form $(\hat{\bm{e}}^{\prime}_{\alpha})\mapsto(0,\hat{\Gamma}^{\prime})$ on $\mathcal{F}_{\hat{\bm{G}_{0}}}$ such that
@@
    The section of the fiber bundle $\pi:\hat{\mathcal{U}}\rightarrow\mathcal{U}$ is an embedding of $\mathcal{U}$ into $\hat{\mathcal{U}}$. The subspace $H_{f(\bm{X})}=(Tf)T_{\bm{X}}\mathcal{U}$ of $T_{f(\bm{X})}\hat{\mathcal{U}}$ is isomorphic to $T_{\bm{X}}\mathcal{U}$ then of dimension 4 and the tangent space to the fiber $\hat{\mathcal{U}}_{\bm{X}}$ at $f(\bm{X})$ is the kernel of $T\pi$. Then $T_{f(\bm{X})}\hat{\mathcal{U}}=H_{f(\bm{X})}\oplus T_{f(\bm{X})}\hat{\mathcal{U}}_{\bm{X}}$. A basis adapted to the section $f$ is a fibered basis $(\hat{\bm{e}}_{1},\cdot,\hat{\bm{e}}_{5})$ such that $(\hat{\bm{e}}_{1},\cdot,\hat{\bm{e}}_{4})$ is a basis of $H_{f(\bm{X})}$. A way to build such a basis is to choose a non null vector $\hat{\bm{e}}_{5}$ tangent to the fiber $\hat{\mathcal{U}}_{\bm{X}}$ and a basis $(\bm{e}_{1},\cdot,\bm{e}_{4})$ of $T_{\bm{X}}\mathcal{U}$. Then, with the convention that Greek indices run from 1 to 5 and Latin ones run from 1 to 4, $(\hat{\bm{e}}_{1},\cdot,\hat{\bm{e}}_{4},\hat{\bm{e}}_{5})=((Tf)\bm{e}_{1},\cdot,(Tf)\bm{e}_{4},\hat{\bm{e}}_{5})$ is a basis adapted to the section $f$ and
@@
    In terms of the dual bases, one has